OpenStudy (anonymous):

Olivia drives her car at 60 miles per hour for t hours. Ian drives his car at 50 miles per hour for (t + 2) hours. Write a simplified expression for the sum of the distances traveled by the two cars.

OpenStudy (saifoo.khan):

Let d = sum of distance travelled by both cars d = 60t + 50(t + 2) d = 60t + 50t + 100 d = 110t + 100

OpenStudy (lalaly):

d = 60t + 50t + 100 d = 110t + 100

OpenStudy (anonymous):

d = 60t + 50t + 100 d = 110t + 100